# FFAA: Multimodal Large Language Model based Explainable Open-World Face Forgery Analysis Assistant

### v1

@@ Abstract
The rapid advancement of deepfake technologies has sparked widespread public concern, particularly as face forgery poses a serious threat to public information security. However, the unknown and diverse forgery techniques, varied facial features and complex environmental factors pose significant challenges for face forgery analysis. Existing datasets lack descriptions of these aspects, making it difficult for models to distinguish between real and forged faces using only visual information amid various confounding factors. In addition, existing methods do not yield user-friendly and explainable results, complicating the understanding of the model’s decision-making process. To address these challenges, we introduce a novel Open-World Face Forgery Analysis VQA (OW-FFA-VQA) task and the corresponding benchmark. To tackle this task, we first establish a dataset featuring a diverse collection of real and forged face images with essential descriptions and reliable forgery reasoning. Base on this dataset, we introduce FFAA: Face Forgery Analysis Assistant, consisting of a fine-tuned Multimodal Large Language Model (MLLM) and Multi-answer Intelligent Decision System (MIDS). By integrating hypothetical prompts with MIDS, the impact of fuzzy classification boundaries is effectively mitigated, enhancing the model’s robustness. Extensive experiments demonstrate that our method not only provides user-friendly explainable results but also significantly boosts accuracy and robustness compared to previous methods.
Introduction
The widespread availability of facial data online and access to deepfake technologies (FaceSwapDevs 2019; Kowalski 2018) have facilitated the forgery of identities. Malicious actors employ deepfake technology to manipulate faces for fraudulent purposes, generating false information and inciting public panic, which pose serious threats to both personal and public information security. To combat face forgery attacks, researchers have analyzed a variety of face forgery datasets (Rossler et al. 2019; Li et al. 2020) and developed numerous advanced forgery detection models (Dong et al. 2023; Ba et al. 2024) utilizing deep learning techniques.
Figure 1: Architecture of FFAA (Left). FFAA achieves state-of-the-art generalization performance on the OW-FFA-Bench and demonstrates excellent robustness (Right).
@@ Introduction
Motivated by these, we introduce a novel Open-World Face Forgery Analysis VQA (OW-FFA-VQA) task, extending the original binary classification task into a VQA task, aimed at providing a user-friendly and reasonable authenticity analysis of the input face image. To support this task, we employ GPT4-assisted data generation to create the FFA-VQA, an explainable open-world face forgery analysis VQA dataset. This dataset includes a diverse collection of real and forged face images with essential descriptions and reliable forgery reasoning. Based on this dataset, we introduce FFAA: Face Forgery Analysis Assistant. Specifically, to endow the model with comprehensive face forgery analysis capabilities and enable flexible responses based on varying hypotheses, we fine-tune an MLLM on the FFA-VQA dataset containing hypothetical prompts. Subsequently, Multi-answer Intelligent Decision System (MIDS) is proposed to select the answer that best matches the image’s authenticity from MLLM’s responses under different hypotheses, mitigating the impact of fuzzy classification boundaries between real and forged faces, thereby enhancing the model’s robustness.
To evaluate model performance under complex real-world conditions, we establish a challenging Open-World Face Forgery Analysis benchmark (OW-FFA-Bench), with a diverse collection of real and forged face images. We also conduct ablation studies to demonstrate the effectiveness of our dataset and the efficacy of MIDS. Furthermore, we evaluate the interpretability of our method by comparing it with advanced MLLMs and visualizing attention heatmaps. Additionally, we provide examples showing that FFAA can effectively distinguish between easy and difficult classification samples without requiring threshold settings. Our method not only provides user-friendly explainable face forgery analysis but also significantly boosts accuracy and model’s robustness compared to previous methods. In summary, our contributions are as follows:
@@
  • Extensive experiments demonstrate that our method not only provides user-friendly explainable results but also significantly boosts accuracy and robustness compared to previous methods.
@@
Related Work
Face Forgery Analysis Datasets and Methods
FF++ (Rossler et al. 2019) is the first large-scale face forgery video dataset, which includes four types of forgery techniques: DeepFakes (FaceSwapDevs 2019), FaceSwap (Kowalski 2018), Face2Face (Thies et al. 2016) and NeuralTextures (Thies, Zollhöfer, and Nießner 2019). Afterwards, some real-world datasets emerge, such as DFDC (Dolhansky et al. 2020), Deeperforensics (Jiang et al. 2020). Although these datasets providing numerous images or videos, they lack the necessary textual descriptions. Recently, Zhang et al. construct the DD-VQA dataset by manually incorporating reasoning process for certain real and forged faces that could be discerned on common sense. However, numerous face forgeries in open-world scenarios are challenging to identify on common sense and require expert scrutiny.
Despite the datasets, researchers have previously proposed numerous methods (Chollet 2017; Zhao et al. 2021; Cao et al. 2022) for detecting face forgeries. However, these methods exhibit limited generalization in open-world scenarios. To improve generalization, researchers have made improvements from various perspectives. Dong et al. find that inadvertently learned identity representations in images hinder generalization thus design a specialized network to mitigate this impact. Ba et al. design a deep forgery detection framework grounded in information bottleneck theory to extract more forgery clues. Recently, Khan and Dang-Nguyen leverage the powerful zero-shot capabilities of CLIP (Radford et al. 2021), employing Prompt-Tuning with a learnable context to integrate textual and visual information for deepfake detection, demonstrating better generalization.
However, these methods still demonstrate limited generalization in complex open-world scenarios. Additionally, these methods typically approach the problem as a binary classification task, yielding solely binary outputs or heatmaps, making it difficult to understand the model’s decision-making process. Therefore, it is essential to develop an open-world face forgery analysis model with strong generalization and robustness, while also providing user-friendly and explainable results.
Multimodal Large Language Models
Recently, Multimodal Large Language Models (MLLMs) represented by GPT-4V and GPT-4o have garnered significant attention due to their remarkable image understanding and analysis capabilities. Some studies (Jia et al. 2024; Shi et al. 2024) have explored their potential in face forgery analysis by employing prompt engineering. They find that carefully designed prompts are essential for inducing analysis and making judgments, yet the accuracy of discrimination remains suboptimal. In other fields, some studies (Lai et al. 2024; Li et al. 2024) employ instruction tuning to fine-tune pre-trained MLLMs, resulting in domain-specific MLLMs with strong zero-shot capabilities.
Dataset
Data Collection
@@
Fig. 2a shows the data collection process. Based on some studies (Tolosana et al. 2020; Dang et al. 2020) and real-world scenarios, we classify face forgery types into three categories: identity exchange (i.e., replacing the target face with another face), facial attribute manipulation (i.e., manipulating attributes of the target face, including expressions), and entire face synthesis (i.e., generating non-existent faces). Then, we utilize FF++ (Rossler et al. 2019), Celeb-DF-v2 (Li et al. 2020), DFFD (Dang et al. 2020) and GanDiffFace (Melzi et al. 2023) as our source datasets since they encompass diverse forgery techniques and enriched facial characteristics. Finally, we construct the Multi-attack (MA) dataset, comprising 95K images with diverse facial features and various types of forgeries.
GPT4-assisted Analysis Generation
@@
Manually writing the image description and forgery reasoning requires specialized knowledge and is quite cumbersome, making it difficult to establish a large-scale and high-quality dataset. Inspired by some studies (Chen et al. 2023; Liu et al. 2024b; Li et al. 2024) using GPT for image captioning, we employ GPT-4o to generate face forgery analysis processes. The pipeline of analysis generation is shown in Fig. 2b, which can be summarized as follows. First, we integrate a carefully designed system prompt, a data source-specific prompt, the target image and available reference images as inputs to GPT-4o to generate responses in a specific format. Subsequently, responses with incorrect formats, erroneous results or low probability are filtered out. Finally, experts scrutinize the analysis process and the qualified responses are organized into a VQA format.
Prompt Design.
@@
Given the challenges of face forgery analysis in open-world scenarios, it is crucial to supply domain-specific auxiliary information and carefully crafted prompts to guide the generation. Motivated by these, our system prompt comprises four sections: role-playing, the introduction to forgery types, the Chain of Thought and conversation requirements. To leverage relevant domain knowledge of GPT-4o and generate responses with domain-specific characteristics, we first introduce its role and the task. Then, the types of face forgeries are defined to avoid ambiguity and assist in the analysis. For the Chain of Thought, we advocate that conducting complex face forgery analysis demands a robust foundation in image understanding, requiring multi-perspective analysis before delivering the final discrimination result. Thus, we design a progressive Face Forgery Analysis Chain of Thought (FFA-CoT) consisting of ’Image description’, ’Forgery reasoning’ and ’Analysis result’ (including classification result, probability and forgery type). Finally, some requirements are claimed to ensure the consistency of the responses. To enhance the reliability of the response, the authenticity and the forgery type of the image are indicated in the data source specific prompt. Additionally, we provide the original face image or forged mask as reference to assist GPT-4o in the analysis.
@@
Quality Verification.
To ensure data quality, responses with incorrect formats, erroneous results or low probability are first filtered out. Subsequently, the images, reference images and responses are provided to experts for scrutinizing the analysis process. Finally, we obtain 20K high-quality face forgery analysis data which are then supplemented with semantically identical inquiry prompts to form the FFA-VQA dataset. In addition, we conduct experiments to verify the quality of FFA-VQA. As depicted in Fig. 3, incorporating high-quality image descriptions and forgery reasoning enables MLLMs to perform face forgery analysis based on image understanding before making final judgments, which effectively reduces overfitting and enhances generalization. See Ablation Study for more experimental details.
@@
Method
@@
Although fine-tuning MLLMs on the FFA-VQA effectively enhances generalization, the model’s robustness remains suboptimal. We also observe that prompts with similar semantics but varied content can yield different authenticity judgments for faces that are challenging to discern. This variability arises from the advancement of forgery techniques and the complexities of open-world scenarios, which have considerably blurred the boundaries between real and forged faces. Motivated by these, we introduce hypothetical prompts to the FFA-VQA dataset that presume the face is either real or fake prior to analysis. By fine-tuning the MLLM on this dataset, we enable the generation of answers based on varying hypotheses. Subsequently, we propose MIDS to select the answer that best matches the image’s authenticity, which effectively mitigates the impact of fuzzy classification boundaries between real and forged faces.
Face Forgery Analysis Assistant
Fig. 4a depicts the architecture of FFAA, which primarily consists of two modules: a fine-tuned MLLM and MIDS. Initially, the MLLM generates the anchor answer $\bm{X}_{a}^{a}$ without hypotheses, the positive answer $\bm{X}_{a}^{p}$ and the negative answer ${\bm{X}_{a}^{p}}$ based on varying hypotheses, all of which are input into MIDS along with the face image. The model then calculates the match score between the image and each answer based on MIDS’s output vectors $(\bm{m}^{a},\bm{m}^{p},\bm{m}^{n})$ and the determination result of each answer as follows:
@@
$s=\begin{cases}m_{0}/(m_{0}+m_{2}),&\text{if }r(\bm{X}_{a})~{}~{}\text{is}~{}~{}^{\prime}real^{\prime},\\ m_{3}/(m_{3}+m_{1}),&\text{if }r(\bm{X}_{a})~{}~{}\text{is}~{}~{}^{\prime}fake^{\prime},\end{cases}$ (1)
@@
where $m_{i}$ represents the $i$-th element of $\bm{m}$ and $r(\bm{X}_{a})$ represents the determination result of $\bm{X}_{a}$. Finally, the answer with the highest match score is selected as the best answer.
Fine-tuning MLLM with Hypothetical Prompts
@@
We categorize the prompts into two types: non-hypothetical (e.g., The image is a human face image. Is it real or fake? Why?) and hypothetical (e.g., This is a [real/fake] human face. What evidence do you have?). We randomly select one-third of the conversations from FFA-VQA and modify their prompts from non-hypothetical to hypothetical, aligning the hypothesis with the authenticity of the face. Then, we fine-tune an MLLM on this dataset with LoRA (Hu et al. 2021). For the image $\bm{X}_{v}$, its conversation data is $(\tilde{\bm{X}}_{q},\bm{X}_{a})$, where $\tilde{\bm{X}}_{q}=\bm{X}_{q}$ for non-hypothetical prompts and $\tilde{\bm{X}}_{q}=[\bm{X}_{h},\bm{X}_{q}]$ for hypothetical prompts. The fine-tuning objective is consistent with the original auto-regressive training objective of the LLM. For a sequence of length $L$, the probability of the target answer is computed as:
$p(\bm{X}_{a}|\bm{X}_{v},\tilde{\bm{X}}_{q})=\prod_{i=1}^{L}p_{\bm{\theta}}(x_{i}|\bm{X}_{v},\tilde{\bm{X}}_{q},\bm{X}_{a,<i}),$ (2)
@@
where $\bm{\theta}$ is the trainable parameters and $\bm{X}_{a,<i}$ are answer tokens before current prediction token $x_{i}$. From the conditionals in Eq. 2, it is evident that for prompts with hypothesis, $\bm{X}_{a}$ depends on both $\bm{X}_{h}$ and $\bm{X}_{v}$. For faces that are challenging to discern, where altering $\bm{X}_{h}$ may result in varying authenticity judgments, we label them as ’Hard’. For easily distinguishable faces, even if altered, the model is likely to rely on the high-weight image tokens in $\bm{X}_{v}$ to generate $\bm{X}_{a}$, maintaining its original judgment. We label these as ’Easy’. Therefore, by simply altering $\bm{X}_{h}$, we can identify faces that are challenging to distinguish and subsequently make a more informed decision among multiple answers.
Multi-answer Intelligent Decision System

### v2

@@ Abstract
The rapid advancement of deepfake technologies has sparked widespread public concern, particularly as face forgery poses a serious threat to public information security. However, the unknown and diverse forgery techniques, varied facial features and complex environmental factors pose significant challenges for face forgery analysis. Existing datasets lack descriptive annotations of these aspects, making it difficult for models to distinguish between real and forged faces using only visual information amid various confounding factors. In addition, existing methods fail to yield user-friendly and explainable results, hindering the understanding of the model’s decision-making process. To address these challenges, we introduce a novel Open-World Face Forgery Analysis VQA (OW-FFA-VQA) task and its corresponding benchmark. To tackle this task, we first establish a dataset featuring a diverse collection of real and forged face images with essential descriptions and reliable forgery reasoning. Based on this dataset, we introduce FFAA: Face Forgery Analysis Assistant, consisting of a fine-tuned Multimodal Large Language Model (MLLM) and Multi-answer Intelligent Decision System (MIDS). By integrating hypothetical prompts with MIDS, the impact of fuzzy classification boundaries is effectively mitigated, enhancing model robustness. Extensive experiments demonstrate that our method not only provides user-friendly and explainable results but also significantly boosts accuracy and robustness compared to previous methods.
1 Introduction
The widespread availability of facial data online and access to deepfake technologies [10, 19] have facilitated the forgery of identities. Malicious actors employ deepfake technologies to manipulate faces for fraudulent purposes, generating false information and inciting public panic, which pose serious threats to both personal and public information security. To combat face forgery attacks, researchers have analyzed a variety of face forgery datasets [36, 22, 8] and developed numerous advanced forgery detection models [9, 3, 33, 4, 39] utilizing deep learning techniques.
Figure 1: Left: Architecture of FFAA. Right: FFAA achieves state-of-the-art generalization performance on OW-FFA-Bench (ACC=86.5%) and exhibits excellent robustness (sACC=10.0%).
@@ 1 Introduction
Motivated by these, we introduce a novel Open-World Face Forgery Analysis VQA (OW-FFA-VQA) task, extending the original binary classification task into a VQA task that, given any face image, requires authenticity determination and an analysis process with supporting evidence. To support this task, we employ GPT4-assisted data generation to create the FFA-VQA, an explainable open-world face forgery analysis VQA dataset, which includes a diverse collection of real and forged face images with essential descriptions and reliable forgery reasoning. Based on this dataset, we introduce FFAA: Face Forgery Analysis Assistant. Specifically, to endow the model with comprehensive face forgery analysis capabilities and enable flexible responses based on varying hypotheses, we fine-tune an MLLM on the FFA-VQA dataset containing hypothetical prompts. Subsequently, Multi-answer Intelligent Decision System (MIDS) is proposed to select the answer that best matches the image’s authenticity from MLLM’s responses under different hypotheses, mitigating the impact of fuzzy classification boundaries between real and forged faces, thereby enhancing model robustness.
To evaluate model performance in complex real-world conditions, we establish a challenging Open-World Face Forgery Analysis benchmark (OW-FFA-Bench), comprising a diverse set of real and forged face images from seven public datasets. Ablation studies are conducted to validate the effectiveness of our dataset and MIDS’s efficacy. Furthermore, we assess the interpretability of our method by comparing it with advanced MLLMs and visualizing attention heatmaps. Examples are also provided to demonstrate FFAA’s capability to distinguish between easy and challenging classification samples without threshold adjustments. Our method not only provides user-friendly, explainable results but also achieves notable improvements in accuracy and robustness over previous methods. To our knowledge, we are the first to explore and effectively utilize fine-tuned MLLMs for explainable face forgery analysis. In summary, our contributions are as follows:
@@
  • Extensive experiments demonstrate that our method significantly enhances accuracy and robustness while providing explainable results compared to previous methods.
2 Related Work
Face Forgery Analysis Datasets and Methods: FF++ [36] is the first large-scale face forgery video dataset, which includes four types of forgery techniques: DeepFakes [10], FaceSwap [19], Face2Face [43] and NeuralTextures [44]. Afterwards, some real-world datasets emerge, such as DFDC [8], Deeperforensics [15]. Although these datasets providing numerous images or videos, they lack the necessary textual descriptions. Recently, Zhang et al. construct the DD-VQA dataset by manually incorporating reasoning process for certain real and forged faces discernible by common sense. However, manual annotation is quite cumbersome and prone to bias. Additionally, numerous face forgeries in open-world scenarios are challenging to identify using common sense alone and require expert scrutiny.
Despite the datasets, researchers have previously proposed numerous methods [6, 49, 4, 29, 39] for detecting face forgeries. However, these methods exhibit limited generalization in open-world scenarios. To improve generalization, researchers have made improvements from various perspectives. Dong et al. find that inadvertently learned identity representations in images hinder generalization thus design a specialized network to mitigate this impact. Lin et al. leverages annotated demographic information [47] and employs disentanglement learning to extract domain-agnostic forgery features and demographic features, fostering fair learning. However, these methods still demonstrate limited generalization in complex open-world scenarios. Additionally, these methods typically approach the problem as a binary classification task, yielding solely binary outputs or heatmaps, making it difficult to understand the model’s decision-making process. Therefore, it is essential to develop an open-world face forgery analysis model with strong generalization and robustness, while also providing user-friendly and explainable results.
Multimodal Large Language Models: Recently, Multimodal Large Language Models (MLLMs) represented by GPT-4V [2] and GPT-4o [32] have garnered significant attention due to their remarkable image understanding and analysis capabilities. Some studies [14, 38] have explored their potential in face forgery analysis by employing prompt engineering, finding that well-crafted prompts are crucial for analysis and judgment, though discrimination accuracy remains suboptimal. In other fields, some studies [20, 21] employ instruction tuning to fine-tune pre-trained MLLMs, resulting in domain-specific MLLMs with strong zero-shot capabilities. However, the fine-tuning of MLLMs for face forgery analysis and the strategies to leverage their advantages remain underexplored.
3 Dataset
@@
3.1 Data Collection
@@
Fig. 2a shows the data collection process. Based on some studies [45, 7] and real-world scenarios, we classify face forgery types into three categories: identity exchange (i.e., replacing the target face with another face), facial attribute manipulation (i.e., manipulating attributes of the target face, including expressions), and entire face synthesis (i.e., generating non-existent faces). Then, we utilize FF++ [36], Celeb-DF-v2 [22], DFFD [7] and GanDiffFace [30] as our source datasets since they encompass diverse forgery techniques and enriched facial characteristics. Finally, we create the Multi-attack (MA) dataset, consisting of 95K images with diverse facial features and multiple forgery types.
3.2 GPT4-assisted Analysis Generation
@@
Manually writing the image description and forgery reasoning requires specialized knowledge and is quite cumbersome, making it difficult to establish a large-scale and high-quality dataset. Inspired by some studies [5, 27, 21] using GPT for image captioning, we employ GPT-4o [32] to generate face forgery analysis. The pipeline of analysis generation is shown in Fig. 2b, which can be summarized as follows. First, a system prompt, forgery-specific prompt, target image and available reference images are integrated as inputs to GPT-4o to produce responses in a structured format. Responses with incorrect formats, erroneous results, or low probability are filtered out. Finally, experts scrutinize the analysis, and qualified responses are organized into a VQA format. See Supplementary Material for more details.
@@
System Prompt Design: Given the challenges of face forgery analysis in open-world scenarios, it is crucial to supply domain-specific auxiliary information and carefully crafted prompts to guide the generation. Motivated by these, our system prompt comprises four sections: role-playing, the introduction to forgery types, the Chain-of-Thought and conversation requirements. To leverage relevant domain knowledge of GPT-4o and generate responses with domain-specific characteristics, we first introduce its role and the task. Then, the types of face forgeries are defined to avoid ambiguity and assist in the analysis. For the Chain-of-Thought, we advocate that conducting complex face forgery analysis demands a robust foundation in image understanding, requiring multi-perspective analysis before delivering the final discrimination result. Thus, we design a progressive Face Forgery Analysis Chain of Thought (FFA-CoT) consisting of ’Image description’, ’Forgery reasoning’ and ’Analysis result’ (including classification result, probability and forgery type). Finally, some requirements are claimed to ensure the consistency of the responses.
@@
Bias Mitigation Strategies: (1) Annotator bias: We use GPT-4o with a zero-shot in-context learning paradigm for annotation, mitigating annotator subjective bias; (2) Generated analysis bias: To enhance response reliability, we include the authenticity and forgery type of the target image as prior information in the forgery-specific prompt (e.g., ’The image is a fake face whose forgery type is identity exchange. Please analyze…’), along with the original image or forged mask as a reference for more precise analysis. (3) Analysis result bias: Responses not conforming to FFA-CoT are filtered to ensure structural consistency. Then responses with erroneous results are removed. Lastly, lower probability often indicate higher analysis difficulty and reduced reliability. Thus, responses with probability values below a set threshold are excluded; (4) Analysis rationality bias: Experts scrutinize responses with access to the original image, its authenticity and forgery type, and reference images to assess the rationality of the forgery analysis. Responses deemed unreasonable by experts are filtered out to mitigate potential biases in GPT-4o’s analysis.
Finally, we obtain 20K high-quality face forgery analysis data which are then supplemented with semantically identical inquiry prompts to form the FFA-VQA dataset.
Quality Verification: We conduct experiments to verify the quality of FFA-VQA. As depicted in Fig. 3, incorporating high-quality image descriptions and forgery reasoning enables MLLMs to perform face forgery analysis based on image understanding before making final judgments, which effectively reduces overfitting and enhances generalization. See Sec.6 for more experimental details.
4 Method
@@
Although fine-tuning MLLMs on FFA-VQA effectively enhances generalization, model robustness remains suboptimal. We also observe that prompts with similar semantics but varied content can yield different authenticity judgments for faces that are challenging to discern. This variability arises from the advancement of forgery techniques and the complexities of open-world scenarios, which have considerably blurred the classification boundaries between real and forged faces. Motivated by these, we introduce hypothetical prompts into FFA-VQA, presuming the face to be either real or fake prior to analysis. Fine-tuning the MLLM on this dataset enables answer generation based on different hypotheses. Subsequently, we propose MIDS to select the answer that best aligns the face’s authenticity, which effectively mitigates the impact of fuzzy classification boundaries between real and forged faces. Fig. 4a depicts the architecture of FFAA, which primarily consists of two modules: a fine-tuned MLLM and MIDS. Initially, the MLLM generates the anchor answer $\mathbf{X}_{a}^{a}$ without hypotheses, the positive answer $\mathbf{X}_{a}^{p}$ and the negative answer ${\mathbf{X}_{a}^{p}}$ based on varying hypotheses, all of which are input into MIDS along with the face image. The model then calculates the match score $S_{m}$ between the image and each answer based on MIDS’s output vectors $(\mathbf{m}^{a},\mathbf{m}^{p},\mathbf{m}^{n})$ and the determination result of each answer as follows:
@@
$S_{m}=\begin{cases}m_{0}/(m_{0}+m_{2}),&\text{if}~{}~{}\mathcal{R}(\mathbf{X}_{a})~{}~{}\text{is}~{}~{}^{\prime}real^{\prime},\\ m_{3}/(m_{3}+m_{1}),&\text{if}~{}~{}\mathcal{R}(\mathbf{X}_{a})~{}~{}\text{is}~{}~{}^{\prime}fake^{\prime},\end{cases}$ (1)
@@
where $m_{i}$ represents the $i$-th element of $\mathbf{m}$ and $\mathcal{R}(\mathbf{X}_{a})$ represents the determination result of $\mathbf{X}_{a}$. Finally, the answer with the highest match score is selected as the best answer.
4.1 Fine-tuning MLLM with Hypothetical Prompts
@@
We categorize the prompts into two types: non-hypothetical (e.g., The image is a human face image. Is it real or fake? Why?) and hypothetical (e.g., This is a [real/fake] human face. What evidence do you have?). One-third of the conversations from FFA-VQA are randomly selected, with their prompts modified from non-hypothetical to hypothetical, aligning the hypothesis with the authenticity of the face. Then, we fine-tune an MLLM on this dataset with LoRA [13]. For the image $\mathbf{X}_{v}$, its conversation data is $(\tilde{\mathbf{X}}_{q},\mathbf{X}_{a})$, where $\tilde{\mathbf{X}}_{q}=\mathbf{X}_{q}$ for non-hypothetical prompts and $\tilde{\mathbf{X}}_{q}=[\mathbf{X}_{h},\mathbf{X}_{q}]$ for hypothetical prompts. The fine-tuning objective is consistent with the original auto-regressive training objective of the LLM. For a sequence of length $L$, the probability of the target answer is computed as:
$p(\mathbf{X}_{a}|\mathbf{X}_{v},\tilde{\mathbf{X}}_{q})=\prod_{i=1}^{L}p_{\boldsymbol{\theta}}(x_{i}|\mathbf{X}_{v},\tilde{\mathbf{X}}_{q},\mathbf{X}_{a,<i}),$ (2)
@@
where $\boldsymbol{\theta}$ is the trainable parameters and $\mathbf{X}_{a,<i}$ are answer tokens before current prediction token $x_{i}$. From the conditionals in Eq. 2, it is evident that for prompts with hypothesis $\mathbf{X}_{h}$, $\mathbf{X}_{a}$ depends on both $\mathbf{X}_{h}$ and $\mathbf{X}_{v}$. For faces that are challenging to discern, where altering $\mathbf{X}_{h}$ may result in varying authenticity judgments, we label them as ’Hard’. For easily distinguishable faces, even if altered, the model is likely to rely on the high-weight image tokens in $\mathbf{X}_{v}$ to maintain its original judgment. We label these as ’Easy’. Therefore, by simply altering $\mathbf{X}_{h}$, we can identify faces that are challenging to distinguish and subsequently make a more informed decision among multiple answers.
4.2 Multi-answer Intelligent Decision System